OpenStudy (anonymous):

Suppose that x(f(x))^3 + xf(x) = 30, and that f(3)=2. Find f'(3). f'(3) =

OpenStudy (freckles):

can you differentiate the given equation?

OpenStudy (freckles):

the x(f(x))^3+xf(x)=30

OpenStudy (freckles):

\[\frac{d}{dx}(f(x))^3 \text{ can be found by chain rule }\]

OpenStudy (freckles):

do you know how to do that?

OpenStudy (anonymous):

\[3f(x)^2*f'(x)\]

OpenStudy (freckles):

that's good

OpenStudy (freckles):

\[\frac{d}{dx} x(f(x))^3 = (f(x))^3 \frac{d}{dx} x+ x \frac{d}{dx}(f(x))^3\]

OpenStudy (freckles):

and you already found the derivative of (f(x))^3

OpenStudy (freckles):

\[\frac{d}{dx} x(f(x))^3 = (f(x))^3 \frac{d}{dx} x+ x \frac{d}{dx}(f(x))^3 \\ =(f(x))^3 \frac{dx}{dx}+x 3(f(x))^2f'(x)\]

OpenStudy (freckles):

So you do know I just used the product rule right?

OpenStudy (freckles):

You can apply the product rule to also differentiate xf(x)

OpenStudy (freckles):

Anyways once you have differentiate both sides replace all the x's with 3 since you attempting to find f'(3) and are given f(3)

OpenStudy (anonymous):

Okay i got a little lost can i just put in the 3s in here? \[f(x)^3+3xf(x)^2*f'(x)+f(x)+xf'(x)=0\]

OpenStudy (freckles):

yes replace all the little x's you see with 3

OpenStudy (freckles):

\[f(3)^3+3(3)f(3)^2*f'(3)+f(3)+3f'(3)=0\]

OpenStudy (freckles):

and you f(3) is 2

OpenStudy (freckles):

\[2^3+3(3)2^2f'(3)+2+3f'(3)=0\]

OpenStudy (freckles):

solve that equation for f'(3)

OpenStudy (freckles):

you may want to simplify first

OpenStudy (anonymous):

ok i got -11.5 but I think that's wrong.

OpenStudy (freckles):

I think that might be wrong too

OpenStudy (freckles):

I have answer between -1 and 0

OpenStudy (freckles):

and -11.5 doesn't fit in there

OpenStudy (freckles):

but I could have done something wrong too

OpenStudy (freckles):

\[2^3+3(3)2^2f'(3)+2+3f'(3)=0 \\ 8+36f'(3)+2+f'(3)=0\]

OpenStudy (freckles):

combine like terms

OpenStudy (freckles):

oops left off the 3 in front of the f'(3)

OpenStudy (freckles):

\[2^3+3(3)2^2f'(3)+2+3f'(3)=0 \\ 8+36f'(3)+2+3f'(3)=0\]

OpenStudy (freckles):

ok combine like terms

OpenStudy (freckles):

8+2=? 36f'(3)+3f'(3)=?

OpenStudy (anonymous):

oops I see what I did wrong. -10/39 is the answer.

OpenStudy (freckles):

\[10+39f'(3)=0\]

OpenStudy (freckles):

yah!

OpenStudy (anonymous):

thanks so much

OpenStudy (freckles):

np